OpenStudy (anonymous):

The velocity of an object increases from –10 m/s to –15 m/s in 2.0 s. What is the average acceleration of the object? +2.5 m/s2 –2.5 m/s2 +12.5 m/s2 –12.5 m/s2

OpenStudy (rushwr):

\[Acceleration = \frac{ \Delta V }{ \Delta t }\]

OpenStudy (anonymous):

WHAT DOES THAT MEAN?

OpenStudy (anonymous):

v divided by t?

OpenStudy (rushwr):

delta V = final velocity - initial velocity Delta T = time difference.

OpenStudy (anonymous):

O.O

OpenStudy (rushwr):

Did u get it ? Substitute and get the answer !!!

OpenStudy (anonymous):

I dont understand anything

OpenStudy (rushwr):

Do u understand the question ?

OpenStudy (anonymous):

no

OpenStudy (rushwr):

Haven't u learned this lesson at school ?

OpenStudy (anonymous):

I go to online school and its harder to understand

OpenStudy (anonymous):

q.q

OpenStudy (anonymous):

@TorilovesVampires

OpenStudy (anonymous):

@countonme123

OpenStudy (rushwr):

okai Acceleration is the rate of change of velocity. Velocity is the rate of change of speed. We do have equations for acceleration. But if we are asked to find the average acceleration then we use only the below equation \[Average acceleration = \frac{ Final velocity - initial velocity }{ time difference}\]

OpenStudy (rushwr):

Here the final velocity is -15ms^-1 Initial velocity is -10ms^-1 and the time difference is 2s now substitute and see

OpenStudy (anonymous):

how would I do that?

OpenStudy (rushwr):

\[Acceleration = \frac{ -15 - (-10) }{ 2 } = \frac{ -5 }{ 2 } = -2.5ms ^{-1}\]

OpenStudy (anonymous):

thnk u

OpenStudy (rushwr):

no probleem.